# The $r$-matrix structure on the moduli space of framed Higgs pairs

### v3

@@ 1 Introduction
Isospectral flows of $\Phi(z)$ can be obtained by using certain spectral invariants of $\Phi$ as hamiltonian generators; for example, on the Poisson submanifold of matrices with simple poles at $z=x_{1},\dots,x_{N}$, (i.e. the Gaudin model [11]), one uses $H_{j}=\mathop{\mathrm{res}}\limits_{z=x_{j}}\frac{1}{2}\mathrm{Tr}\Phi^{2}(z)\mathrm{d}z$ as commuting Hamiltonians. Separating Darboux . coordinates for these integrable systems were provided in [1].
Another important example is the Calogero and elliptic-Calogero-Moser (eCM) systems; they both have Lax-matrix formulations. These isospectral dynamical systems were generalized by Hitchin in [14] so that the matrix $\Phi$ depends on a point $p$ in an arbitrary smooth Riemann surface $\mathcal{C}$ (of arbitrary genus). To be more precise, in Hitchin’s construction $\Phi$ is a holomorphic differential with values in ${\rm End}(\mathscr{E})$, where $\mathscr{E}$ is a stable vector bundle over $\mathcal{C}$ of rank $n$. This is called a Higgs field and it can be identified with an element of the co-tangent space of the moduli space of vector bundles at the point $\mathscr{E}$. To recover the rational case one has to consider “generalized” Higgs fields (allowing poles) and specialize $\mathcal{C}$ to $\mathbb{P}^{1}$. In the elliptic case the eCM system can also be presented as a special case of Hitchin’s system [9].
In the elliptic (and trigonometric) case the $r$-matrix structure (1.1) was defined in [24, 9]; for the three cases, rational, trigonometric and elliptic, in fact there are families of pencils of compatible structures, as shown in [13]. The question, however, of a proper $r$-matrix structure for Hitchin’s systems over a curve of arbitrary genus $g$ is precisely the goal of this paper. In fact in [15] the authors conclude “[…] we can see that the Poisson geometry of rational surfaces suggests quite strongly that there is no nice Poisson extension of the Sklyanin bracket to arbitrary base curves”. The use of the Tyurin parametrization and a partial description of the rr–matrix was considered in [6] and hence the present paper does not appear in a vacuum. We comment below in Remark 1.2 on the added value of the present paper.
@@ 2 Framed vector bundles and their moduli
Let $\mathcal{C}$ be a smooth Riemann surface of genus $g$ with complex structure. We consider a generic vector bundle $\mathscr{E}$ of rank $n$ and degree $ng$. The moduli space of such objects is locally a smooth manifold of complex dimension $n^{2}(g-1)+1$. The corresponding co-tangent bundle is canonically a complex-symplectic manifold of twice the dimension: if $q_{\alpha}$ is a local set of coordinates and $p^{\alpha}$ the corresponding momenta (coordinates in the cotangent space), the symplectic structure is expressed as
@@ 2.1 The reproducing kernel on the space of framed Higgs fields
Here $\omega_{1},\dots,\omega_{g}$ are holomorphic differentials normalized to have zero periods along $g$ homologically independent cycles mutually non intersecting, and $\omega_{q,\infty}(p)$ is the unique third kind differential with poles at $p=q,\infty$ and residues 1,11,-1,respectively, and with similarly vanishing periods. Note that this expression makes it clear that the Cauchy kernel is single–valued as a meromorphic function of $q$; analytic continuation along the $b_{\ell}$ cycle adds to the last column one of the previous columns. In particular it is independent of the choice of cycles used to normalize the differentials.
@@
The formula (2.28) may seem to be computationally transcendental because it implies that one has to compute the normalized differentials and hence the matrix of periods along a symplectic basis in homology. This is actually not the case and we provide an alternative choice of basis which is purely algebraic (i.e. does not involve computation of periods) in App. B. $\blacktriangle$
@@
where $\kappa_{j}$ are arbitrary parameters, as well as the matrices $N^{(j)}$ in the manifold of nilpotent, rank-one matrices.
@@ 3 The higher genus $r$-matrix structure
$\left\{\displaystyle\mathop{\Phi}^{1}(w)\displaystyle\mathop{,}^{\otimes}\displaystyle\mathop{\mathbf{K}}^{2}(q,r)\right\}\$$=+\displaystyle\mathop{\mathbf{K}}^{1}(w,r)\displaystyle\mathop{\mathbf{K}}^{2}(q,w)\Pi+\displaystyle\mathop{\mathbf{K}}^{1}(w,q)\displaystyle\mathop{\mathbf{K}}^{1}(q,r)\Pi-\displaystyle\mathop{\mathbf{K}}^{1}(w,r)\displaystyle\mathop{\mathbf{K}}^{2}(q,r)\Pi,$ (3.6)
@@ 4 Example: the elliptic case revisited
The r-matrix structure on an elliptic curve was derived in [24, 9]. To be as concrete as possible we realize the elliptic curve as $\mathcal{E}_{\tau}={\mathbb{C}}/{\mathbb{Z}}+\tau{\mathbb{Z}}$, using the symbols $p,q,$... as global coordinates rather than abstract points.
@@ 4.1 Elliptic Calogero-Moser leaf
It is also appropriate here to comment on the now venerable construction of Krichever [17] of the Lax matrix for the elliptic Calogero-Moser (eCM) systems [^2] Krichever used the Weierstrass functions for the construction while here we use the Jacobi Theta functions, but it is a minor difference; we are here aligned to the presentation of Takasaki [25].. The eCM system is the dynamical system with Hamiltonian
@@
The choice of $C_{ab}=i\gamma$ can be explained in a more general way by reminding of the formulation of rational Calogero system and the Kazhdan-Kostant-Sternberg reduction scheme [16]. Indeed in that case one starts with the symplectic manifold of pairs of hermitian matrices matrices $\mathbb{P},\mathbb{Q}$ with the symplectic structure $\{\mathbb{P}_{ab},\mathbb{Q}_{cd}\}=\delta_{ad}\delta_{cb}$. This is simply the canonical symplectic structure on the co-tangent bundle of the vector space of Hermitian matrices. Then consider the trivial Hamiltonian $H=\frac{1}{2}\mathrm{Tr}\mathbb{P}^{2}$ generating the linear flow, $\mathbb{Q}(t)=t\mathbb{P}+\mathbb{Q}$. The angular momentum $\mathbb{J}:=[\mathbb{P},\mathbb{Q}]$ is conserved. The Calogero leaf here is the choice of angular momentum [16]
@@ A.1 Poisson brackets of the kernel with $\Phi$
Appendix B Alternative construction of the Cauchy kernel

### v4

@@ 1 Introduction
Isospectral flows of $\Phi(z)$ can be obtained by using certain spectral invariants of $\Phi$ as hamiltonian generators; for example, on the Poisson submanifold of matrices with simple poles at $z=x_{1},\dots,x_{N}$, (i.e. the Gaudin model [11]), one uses $H_{j}=\mathop{\mathrm{res}}\limits_{z=x_{j}}\frac{1}{2}\mathrm{Tr}\Phi^{2}(z)\mathrm{d}z$ as commuting Hamiltonians. Separating Darboux coordinates for these integrable systems were provided in [1].
Another important example is the Calogero and elliptic-Calogero-Moser (eCM) systems; they both have Lax-matrix formulations. These isospectral dynamical systems were generalized by Hitchin in [14] so that the matrix $\Phi$ depends on a point $p$ belonging to an arbitrary smooth Riemann surface $\mathcal{C}$ (of arbitrary genus). To be more precise, in Hitchin’s construction $\Phi$ is a holomorphic differential with values in ${\rm End}(\mathscr{E})$, where $\mathscr{E}$ is a stable vector bundle over $\mathcal{C}$ of rank $n$. This is called a Higgs field and it can be identified with an element of the co-tangent space of the moduli space of vector bundles at the point $\mathscr{E}$. To recover the rational case one has to consider “generalized” Higgs fields (allowing poles) and specialize $\mathcal{C}$ to $\mathbb{P}^{1}$. In the elliptic case the eCM system can also be presented as a special case of Hitchin’s system [9].
In the elliptic (and trigonometric) case the $r$-matrix structure (1.1) was defined in [5, 24, 9]; for the three cases, rational, trigonometric and elliptic, in fact there are families of pencils of compatible structures, as shown in [13]. The question, however, of a proper $r$-matrix structure for Hitchin’s systems over a curve of arbitrary genus $g$ is precisely the goal of this paper. In fact in [15] the authors conclude “[…] we can see that the Poisson geometry of rational surfaces suggests quite strongly that there is no nice Poisson extension of the Sklyanin bracket to arbitrary base curves”. Obstruction of similar nature was observed much earlier in [28]. The use of the Tyurin parametrization and a partial description of the rr–matrix was considered in [6] and hence the present paper does not appear in a vacuum. We comment below in Remark 1.2 on the added value of the present paper.
@@ 2 Framed vector bundles and their moduli
Let $\mathcal{C}$ be a smooth Riemann surface of genus $g$ with complex structure. We consider a generic holomorphic vector bundle $\mathscr{E}$ of rank $n$ and degree $ng$. The moduli space of such objects is locally a smooth complex manifold of dimension $n^{2}(g-1)+1$. The corresponding co-tangent bundle is canonically a complex-symplectic manifold of twice the dimension: if $q_{\alpha}$ is a local set of coordinates and $p^{\alpha}$ the corresponding momenta (coordinates in the cotangent space), the symplectic structure is expressed as
@@ 2.1 The reproducing kernel on the space of framed Higgs fields
Here $\omega_{1},\dots,\omega_{g}$ are holomorphic differentials normalized along $g$ homologically independent cycles, mutually non intersecting, $\oint_{a_{j}}\omega_{k}=\delta_{jk}$, and $\omega_{q,\infty}(p)$ is the unique third kind differential with poles at $p=q,\infty$ and residues 1,11,-1, respectively, and with vanishing periods along the same cycles. Note that this expression makes it clear that the Cauchy kernel is single–valued as a meromorphic function of $q$; analytic continuation along the $b_{\ell}$ cycle adds to the last column one of the previous columns. In particular it is independent of the choice of cycles used to normalize the differentials.
@@
The formula (2.28) may seem to be computationally transcendental because it requires to compute the normalized differentials and hence the matrix of periods along a symplectic basis in homology. This is actually not the case and we provide an alternative choice of basis which is purely algebraic (i.e. does not involve computation of periods) in App. B. $\blacktriangle$
@@
(see the notational convention above for the meaning of $q_{j}$) where $\kappa_{j}$ are arbitrary parameters, as well as the matrices $N^{(j)}$ in the manifold of nilpotent, rank-one matrices.
@@
Mittag-Leffler problems.
The Cauchy kernel $\mathbf{K}$ solves two problems of Mittag-Leffler type: suppose we are looking for a meromorphic section, $\sigma$, of $\mathscr{E}$ with a prescribed singular part at a point $r_{0}$:
$\sigma(q)=\sum_{\ell=0}^{L}\sigma_{\ell}\frac{1}{z^{\ell}}+\mathcal{O}(|z|)=\sigma_{sing}(z)+\mathcal{O}(1),$
here $z=z(q)$ is a local coordinate centered at $r_{0}$ ($z(r_{0})=0$). Then the Cauchy kernel allows to turn this local information into a global one by means of
$\sigma(q)=\mathop{\mathrm{res}}\limits_{p=r_{0}}\sigma_{sing}(z(p))\mathbf{K}(p,q),$ (2.33)
providing the unique section with the preassigned singular expansion and also vanishing at $\infty$. On the dual side, if we are looking for a meromorphi section $\phi$ of $\mathscr{E}^{\vee}\otimes\mathcal{K}_{\mathcal{C}}$ then similar task is accomplished with the aid of the formula
$\phi(p)=\mathop{\mathrm{res}}\limits_{q=r_{0}}\mathbf{K}(p,q)\phi_{sing}(z(q)\mathrm{d}z(q).$
@@ 3 The higher genus $r$-matrix structure
$\left\{\displaystyle\mathop{\Phi}^{1}(w)\displaystyle\mathop{,}^{\otimes}\displaystyle\mathop{\mathbf{K}}^{2}(q,r)\right\}\$$=\displaystyle\mathop{\mathbf{K}}^{1}(w,r)\displaystyle\mathop{\mathbf{K}}^{2}(q,w)\Pi+\displaystyle\mathop{\mathbf{K}}^{1}(w,q)\displaystyle\mathop{\mathbf{K}}^{1}(q,r)\Pi-\displaystyle\mathop{\mathbf{K}}^{1}(w,r)\displaystyle\mathop{\mathbf{K}}^{2}(q,r)\Pi,$ (3.6)
@@ 4 Example: the elliptic case revisited
The r-matrix structure on an elliptic curve was derived in [5, 24, 9]. To be as concrete as possible we realize the elliptic curve as $\mathcal{E}_{\tau}={\mathbb{C}}/{\mathbb{Z}}+\tau{\mathbb{Z}}$, using the symbols $p,q,$... as global coordinates rather than abstract points.
@@ 4.1 Elliptic Calogero-Moser leaf
It is also appropriate here to comment on the now venerable construction of Krichever [17] of the Lax matrix for the elliptic Calogero-Moser (eCM) systems [^2] Krichever used the Weierstrass functions for the construction while here we use the Jacobi Theta functions, but it is a minor difference; we are here aligned with the presentation of Takasaki [25].. The eCM system is the dynamical system with Hamiltonian
@@
The choice of $C_{ab}=i\gamma$ can be explained in a more general way by reminding of the formulation of rational Calogero system and the Kazhdan-Kostant-Sternberg reduction scheme [16]. Indeed in that case one starts with the symplectic manifold of pairs of hermitian matrices $\mathbb{P},\mathbb{Q}$ with the symplectic structure $\{\mathbb{P}_{ab},\mathbb{Q}_{cd}\}=\delta_{ad}\delta_{cb}$. This is simply the canonical symplectic structure on the co-tangent bundle of the vector space of Hermitian matrices. Then consider the trivial Hamiltonian $H=\frac{1}{2}\mathrm{Tr}\mathbb{P}^{2}$ generating the linear flow, $\mathbb{Q}(t)=t\mathbb{P}+\mathbb{Q}$. The angular momentum $\mathbb{J}:=[\mathbb{P},\mathbb{Q}]$ is conserved. The Calogero leaf here is the choice of angular momentum [16]
@@ 5 Concluding remarks
Data availability statement. This work did not rely on any additional data.
@@ A.1 Poisson brackets of the kernel with $\Phi$
Appendix B Purely algebraic construction of the Cauchy kernel